OpenStudy (boxman61):

Rationalize the denominator √(5) / √(7)-3

OpenStudy (boxman61):

\[(\sqrt{5}/\sqrt{7}-3)(\sqrt{7}+3/\sqrt{7}+3)\]\[\sqrt{35}+3\sqrt{5}/\sqrt{49}-9\]\[\sqrt{35}3\sqrt{5}/\sqrt{49}-9\]\[\sqrt{35}+3\sqrt{5}/7-9\]\[\sqrt{35}+3\sqrt{5}/-2\] thats where i got stuck